# MAPS: A Multilingual Benchmark for Global Agent Performance and Security

### v1

@@ Abstract
Agentic AI systems, which build on Large Language Models (LLMs) and interact with tools and memory, have rapidly advanced in capability and scope. Yet, since LLMs have been shown to struggle in multilingual settings, typically resulting in lower performance and reduced safety, agentic systems risk inheriting these limitations. This raises concerns about the global accessibility of such systems, as users interacting in languages other than English may encounter unreliable or security-critical agent behavior. Despite growing interest in evaluating agentic AI, existing benchmarks focus exclusively on English, leaving multilingual settings unexplored. To address this gap, we propose MAPS, a multilingual benchmark suite designed to evaluate agentic AI systems across diverse languages and tasks. MAPS builds on four widely used agentic benchmarks — GAIA (real-world tasks), SWE-bench (code generation), MATH (mathematical reasoning), and the Agent Security Benchmark (security). We translate each dataset into ten diverse languages, resulting in 805 unique tasks and 8,855 total language-specific instances. Our benchmark suite enables a systematic analysis of how multilingual contexts affect agent performance and robustness. Empirically, we observe consistent degradation in both performance and security when transitioning from English to other languages, with severity varying by task and correlating with the amount of translated input. Building on these findings, we provide actionable recommendations to guide agentic AI systems development and assessment under multilingual settings. This work establishes a standardized evaluation framework, encouraging future research towards equitable, reliable, and globally accessible agentic AI. MAPS benchmark suite is publicly available.  [^1] https://huggingface.co/datasets/Fujitsu-FRE/Multilingual-Agentic-AI-Benchmark
@@ 1 Introduction
LLM-based agentic AI systems combine multi-step reasoning with external tools and memory to solve open-ended tasks such as code generation, web navigation, planning, and transactional services like booking and ordering (Acharya et al., 2025). By doing so, they extend to complex, real-world problems beyond standard LLM benchmarks. Since such real-world applications serve speakers of diverse languages, maintaining consistent reliability in every language becomes critical. However, since agentic behavior is grounded in LLMs, which often perform inconsistently across languages (Deng et al., 2023; Wang et al., 2023), agents may inherit these multilingual limitations as well, affecting their functionality and trustworthiness. This presents a barrier to equitable access, as non-English users may face degraded responses, incorrect tool actions, or unsafe behaviors—failures that can lead to actual harm in the real world, including erroneous transactions, data corruption, and other security vulnerabilities (Zhang et al., 2024).
@@
To assess emerging agentic systems, various benchmarks have been proposed to evaluate agent performance across a range of tasks (Mialon et al., 2023; Jimenez et al., 2023; Chang et al., 2024; Xu et al., 2024). However, these benchmarks remain English-only. In contrast to multilingual LLM benchmarks (Dang et al., 2024; Shi et al., 2022; Goyal et al., 2022), no equivalent exists for agentic AI tasks—creating a blind spot in our understanding of cross-language performance, safety, and security.
Figure 1: MAPS benchmark suite evaluates LLM-based agents across $11$ languages and $4$ agentic benchmarks covering performance and security.
@@
In this paper, we address this gap. We hypothesize that multilingual settings will reveal performance and security gaps in agentic systems that are not captured by the existing, English-only benchmarks. To investigate this, we introduce MAPS, a Multilingual Agentic AI Benchmark Suite for Performance and Security. MAPS is based on four established agentic benchmarks: GAIA (real-world tasks) (Mialon et al., 2023), SWE-bench (code generation) (Jimenez et al., 2023), MATH (mathematical reasoning) (Hendrycks et al., 2021), and the Agent Security Benchmark (security) (Zhang et al., 2024). These benchmarks are extended to ten typologically diverse languages beyond English [^2] MAPS offers evaluation in the following ten languages: German, Spanish, Portuguese (Brazil), Japanese, Russian, Italian, Arabic, Hebrew, Korean, and Hindi., by employing a hybrid machine and LLM-based translation approach Ki and Carpuat (2024) with extended verification and enhancements. In total, MAPS includes 805 unique tasks, each available in 11 language versions—including the original English and 10 translated variants—for a total of 8,855 instances. An overview of the benchmark structure is shown in Figure 1.
@@
To demonstrate the use of MAPS and test our hypothesis, we selected a leading open-source agent associated with each of the four original benchmarks and applied it to the corresponding multilingual extension. We observed notable declines in both task performance and security when moving from English to other languages, with the severity of these drops varying by task type and correlating with the proportion of non-English input, suggesting that multilingual performance interventions should be targeted based on input composition and task sensitivity. Beyond overall degradation, our findings reveal that multilingual inputs can amplify agentic vulnerabilities in safety-critical tasks, highlighting the need for multilingual risk assessment. These results empirically support our hypothesis and demonstrate the utility of MAPS as a tool for systematic, multilingual evaluation of agentic AI systems.
The primary contributions of this paper are threefold:
@@
  • To the best of our knowledge, we introduce the first multilingual benchmark suite for agentic AI, extending four widely used benchmarks into ten typologically diverse languages for systematic performance and security assessment.
@@ 2.1 Agentic AI Benchmarks
Evaluation objective: performance-oriented benchmarks measure task completion, multi-step reasoning, and correct tool invocation (e.g., AgentBoard (Chang et al., 2024)), whereas security-focused suites probe robustness to adversarial inputs, jailbreaks, and unsafe behaviors (e.g., AgentHarm (Andriushchenko et al., 2024)). Agentic task scope: full-agentic evaluations present only problem statements and expected outcomes, requiring end-to-end planning and execution (e.g., GAIA (Mialon et al., 2023)), while semi-agentic frameworks supply scaffolding, such as code templates or mock APIs, to isolate the LLM’s reasoning and tool-selection core (e.g., AppWorld (Trivedi et al., 2024). Design and evaluation characteristics: most benchmarks span a limited set of domains (three to five use cases), typically including real-world information retrieval and navigation (e.g., AssistantBench (Yoran et al., 2024)), code generation (e.g., SWE-Bench (Jimenez et al., 2023)), reasoning and planning (e.g., MATH (Hendrycks et al., 2021), Travel Planner (Xie et al., 2024)), and security scenarios (e.g., Agent Security Benchmark (Zhang et al., 2024)). They use fixed task counts and predefined difficulty tiers, and to enable reliable, objective measurement despite agents’ open-ended capabilities, they often restrict tasks to closed-form problems with definitive ground truth, allowing clear determination of success or failure (Jimenez et al., 2023; Mialon et al., 2023). A detailed comparison of benchmark design choices, task types, and evaluation properties is provided in the supplementary materials.
@@
While multilingual LLM benchmarks such as XTREME (Hu et al., 2020), FLORES (Goyal et al., 2022), and SIB-200 (Adelani et al., 2023) have enabled broad cross-lingual evaluation, they do not assess interactive decision-making, tool use, or task execution, which are core elements of agentic systems. As a result, existing multilingual benchmarks fall short of capturing the complexities and vulnerabilities that arise when agents operate in non-English settings. This leaves non-English users exposed to agentic failures in their native languages and underscores the need for fully agentic benchmarks that include performance and security evaluation, high data fidelity, and comprehensive multilingual assessment - gaps that our benchmark is specifically designed to address.
2.2 Multilingual Benchmarks and Multilingual Limitations of General-Purpose LLMs
@@
Recent studies show that pre-trained LLMs often struggle with non-English input, especially in languages with limited training resources or those typologically distant from English. Multilingual benchmarks such as XTREME (Hu et al., 2020) and XGLUE(Liang et al., 2020) report consistent accuracy drops when moving from English to languages such as Swahili or Nepali. These gaps reflect an imbalance in pretraining corpora, where English accounts for over $90\%$ of the data, as well as challenges in tokenizing morphologically rich languages and the scarcity of fine-tuning data in many languages (Jha, 2024). Notably, even large models (e.g., GPT-4, Llama 405B) face a “cross-lingual knowledge barrier” on MMLU (Hendrycks et al., 2020) and on Safety tasks (Grattafiori et al., 2024), showing that scale alone does not resolve multilingual performance deficits (Chua et al., 2024).
@@
Building on these performance gaps, LLMs also face robustness and security challenges in multilingual contexts. Since most alignment and red-teaming efforts have been English-centric, models are more prone to generate toxic or policy-violating outputs when processing non-English prompts (Deng et al., 2023; Wang et al., 2023; Aakanksha et al., 2024). Furthermore, hallucination rates increase and confidence calibration degrades outside English, causing models to produce fluent, yet unreliable or potentially harmful content in undersupported languages (Xue et al., 2024). Although security interventions, such as multilingual alignment, have been shown to be effective in reducing harmful output between languages, they often incur a measurable cost in downstream performance or increased latency (Aakanksha et al., 2024).
@@
Given that agentic AI systems are based directly on these LLMs, we hypothesize that they inherit the same language-dependent performance and security limitations. As these agents carry out real-world tasks such as executing code, querying external tools, or navigating web environments, any inherited shortcomings can lead to severe consequences. Yet, to our knowledge, no systematic evaluation has probed how these vulnerabilities manifest within agentic systems. To address this gap, we introduce MAPS, our multilingual agentic benchmark suite in Section 3.

### v2

# MAPS: A Multilingual Benchmark for Agent Performance and Security

@@ Abstract
Agentic AI systems, which build on Large Language Models (LLMs) and interact with tools and memory, have rapidly advanced in capability and scope. Yet, since LLMs have been shown to struggle in multilingual settings, typically resulting in lower performance and reduced safety, agentic systems risk inheriting these limitations. This raises concerns about the accessibility of such systems, as users interacting in languages other than English may encounter unreliable or security-critical agent behavior. Despite growing interest in evaluating agentic AI, existing benchmarks focus exclusively on English, leaving multilingual settings unexplored. To address this gap, we propose MAPS, a multilingual benchmark suite designed to evaluate agentic AI systems across diverse languages and tasks. MAPS builds on four widely used agentic benchmarks — GAIA (real-world tasks), SWE-bench (code generation), MATH (mathematical reasoning), and the Agent Security Benchmark (security). We translate each dataset into eleven diverse languages, resulting in 805 unique tasks and 9,660 total language-specific instances - enabling a systematic analysis of the multilingual effect on AI agents’ performance and robustness. Empirically, we observe degradation in both performance and security when transitioning from English to other languages, with severity varying by task and correlating with the amount of translated input. Building on these findings, we provide actionable recommendations to guide agentic AI systems development and assessment under multilingual settings. This work establishes the first standardized evaluation framework for multilingual agentic AI, encouraging future research towards equitable, reliable, and accessible agentic AI. MAPS benchmark suite will become publicly available upon publication.  [^1] https://huggingface.co/datasets/Fujitsu-FRE/MAPS
Abstract
This supplementary material provides additional details supporting our MAPS benchmark, including extended benchmark comparisons, experimental protocols, and multilingual evaluation results. It is organized as follows:
  • Reproducibility Resources – links to the full datasets and codebase, along with licensing terms (MIT) to enable open use and distribution for research and development purposes.
  • Agentic AI Benchmark Comparison – an expanded analysis of existing performance and security agentic benchmarks across dimensions such as evaluation objective, task type, answer format, and multilingual support.
  • Hybrid Translation Pipeline – additional information on our hybrid translation methodology, including domain-specific prompt design and translation configuration experiments.
  • Manual Evaluation of Translation Quality – annotation guidelines, evaluation examples, and a complete breakdown of human verification scores per dataset and language.
  • Experimental Settings – full specification of the technical setup, agent configurations, and language adaptation process used to isolate multilingual effects across tasks.
  • Extended Results – detailed multilingual performance tables for GAIA, MATH, SWE-Bench, and ASB across individual evaluation runs.
MAPS: A Multilingual Benchmark for Agent Performance and Security
Omer Hofman1, Oren Rachmil1, Shamik Bose1, Vikas Pahuja1, Jonathan Brokman1, Toshiya Shimizu2, Trisha Starostina3, Kelly Marchisio3, Seraphina Goldfarb-Tarrant3, Roman Vainshtein1 1Fujitsu Research of Europe  2Fujitsu Limited  3Cohere
Figure 1: MAPS benchmark suite evaluates LLM-based agents across $12$ languages and 44 agentic benchmarks covering performance and security.
@@ 1 Introduction
LLM-based agentic AI systems combine multi-step reasoning with external tools and memory to solve open-ended tasks such as code generation, web navigation, planning, and transactional services like booking and ordering (Acharya et al., 2025). By doing so, they extend to complex, real-world problems beyond standard LLM benchmarks. Since such real-world applications serve speakers of diverse languages, maintaining consistent reliability in every language becomes critical. However, since agentic behavior is grounded in LLMs, which often perform inconsistently across languages (Deng et al., 2023; Wang et al., 2023), agents may inherit these multilingual limitations as well, affecting their functionality and trustworthiness. This presents a barrier to equitable access, as non-English users may face degraded responses, incorrect tool actions, or unsafe behaviors—failures that can lead to actual harm in the real world, including erroneous transactions, data corruption, and security vulnerabilities (Zhang et al., 2024).
@@
To assess emerging agentic systems, various benchmarks have been proposed to evaluate agent performance across a range of tasks (Mialon et al., 2023; Jimenez et al., 2023; Chang et al., 2024; Xu et al., 2024). However, these benchmarks remain English-only. In contrast to multilingual LLM benchmarks (Dang et al., 2024; Shi et al., 2022; Goyal et al., 2022), no equivalent exists for agentic AI—creating a blind spot in our understanding of cross-language performance, safety, and security.
@@
In this paper, we address this gap. We hypothesize that multilingual settings will reveal performance and security gaps in agentic systems that are not captured by the existing, English-only benchmarks. To investigate this, we introduce MAPS, a Multilingual Agentic AI Benchmark Suite for Performance and Security. MAPS is based on four established agentic benchmarks: GAIA (real-world tasks) (Mialon et al., 2023), SWE-bench (code generation) (Jimenez et al., 2023), MATH (mathematical reasoning) (Hendrycks et al., 2021), and the Agent Security Benchmark (security) (Zhang et al., 2024). These benchmarks are extended to eleven typologically diverse languages beyond English [^2] MAPS offers evaluation in twelve languages: English, German, Spanish, Portuguese (Brazil), Japanese, Russian, Italian, Arabic, Hebrew, Korean, Chinese and Hindi., by employing a hybrid machine and LLM-based translation approach Ki and Carpuat (2024) with extended verification and enhancements. In total, MAPS includes 805 unique tasks, each available in 12 language versions—including the original English and 11 translated variants—for a total of 9,660 instances. An overview of the benchmark layout is shown in Fig. 1.
@@
To demonstrate the use of MAPS and test our hypothesis, we selected a leading open-source agent associated with each of the four original benchmarks and applied it to the corresponding multilingual extension. We observed notable declines in both task performance and security when moving from English to other languages, with the severity of these drops varying by task type and correlating with the proportion of non-English input tokens, suggesting that multilingual performance interventions should be targeted based on input composition and task sensitivity. Beyond overall degradation, our findings reveal that multilingual inputs can amplify agentic vulnerabilities in safety-critical tasks, highlighting the need for multilingual risk assessment. These results empirically support our hypothesis and demonstrate the utility of MAPS for multilingual evaluation of agentic AI systems.
This paper makes three primary contributions:
@@
  • To the best of our knowledge, we introduce the first multilingual benchmark suite for agentic AI, extending four widely used benchmarks into eleven typologically diverse languages for systematic performance and security assessment.
@@ 2.1 Agentic AI Benchmarks
Evaluation objective: Performance-oriented benchmarks measure task completion, multi-step reasoning, and correct tool use (e.g., AgentBoard (Chang et al., 2024)), whereas security-focused suites probe robustness to adversarial inputs, jailbreaks, and unsafe behaviors (e.g., AgentHarm (Andriushchenko et al., 2024)). Agentic task scope: Full-agentic evaluations present only problem statements and expected outcomes, requiring end-to-end planning and execution (e.g., GAIA (Mialon et al., 2023)), while semi-agentic frameworks supply scaffolding, such as code templates or mock APIs, to isolate the LLM’s reasoning and tool-selection core (e.g., AppWorld (Trivedi et al., 2024). Design and evaluation characteristics: Most benchmarks span a limited set of domains (three to five use cases), typically including real-world information retrieval and navigation (e.g., AssistantBench (Yoran et al., 2024)), code generation (e.g., SWE-Bench (Jimenez et al., 2023)), reasoning and planning (e.g., MATH (Hendrycks et al., 2021), and security scenarios (e.g., Agent Security Benchmark (Zhang et al., 2024)). They enable objective measurement despite agents’ open-ended capabilities, by formulating tasks with definitive ground truth, allowing clear determination of success (Jimenez et al., 2023; Mialon et al., 2023). Further comparisons are provided in the Appendix.
@@
While multilingual LLM benchmarks such as XTREME (Hu et al., 2020), FLORES (Goyal et al., 2022), and SIB-200 (Adelani et al., 2023) have enabled broad cross-lingual evaluation, they do not assess interactive decision-making, tool use, or task execution - the core elements of agentic systems. Thus, existing multilingual benchmarks fail to capture the complexities and vulnerabilities agents face in non-English settings - underscoring the need for multilingual agentic benchmarks.
2.2 LLMs Multilingual Limitations
@@
Studies show that pre-trained LLMs struggle with non-English inputs, particularly in languages that are low-resource or typologically distant from English. Multilingual benchmarks (Hu et al., 2020; Liang et al., 2020) report accuracy drops when moving from English to languages such as Swahili or Nepali. These gaps stem from imbalanced training data, with $90\%$ English data, as well as challenges in tokenizing morphologically rich languages and limited multilingual fine-tuning data (Jha, 2024). Notably, even large models (e.g., GPT-4, Llama 405B) face a “cross-lingual knowledge barrier” (Hendrycks et al., 2020), (Chua et al., 2024).
@@
LLMs also face security risks in multilingual contexts. Since most security efforts have been English-centric, models are more prone to generate policy-violating outputs when processing non-English prompts (Deng et al., 2023; Wang et al., 2023; Aakanksha et al., 2024). Multilingual safety alignment, have been shown to be effective but it also degrades general performance latency (Aakanksha et al., 2024).
@@
Since agentic AI systems are based on these LLMs, we hypothesize that they inherit the same language-dependent performance and security limitations. These agents carry out real-world tasks , thus any inherited shortcomings can lead to severe consequences. To our knowledge, no systematic evaluation has examined these vulnerabilities in agentic systems. To address this gap, we introduce MAPS, a multilingual agentic benchmark suite.